# MASSV: Multimodal Adaptation and Self-Data Distillation for Speculative Decoding of Vision-Language Models

### v1

@@ Abstract
Speculative decoding significantly accelerates language model inference by enabling a lightweight draft model to propose multiple tokens that a larger target model verifies simultaneously. However, applying this technique to vision-language models (VLMs) presents two fundamental challenges: small language models that could serve as efficient drafters lack the architectural components to process visual inputs, and their token predictions fail to match those of VLM target models that consider visual context. We introduce Multimodal Adaptation and Self-Data Distillation for Speculative Decoding of Vision-Language Models (MASSV), which transforms existing small language models into effective multimodal drafters through a two-phase approach. MASSV first connects the target VLM’s vision encoder to the draft model via a lightweight trainable projector, then applies self-distilled visual instruction tuning using responses generated by the target VLM to align token predictions. Comprehensive experiments across the Qwen2.5-VL and Gemma3 model families demonstrate that MASSV increases accepted length by up to 30% and delivers end-to-end inference speedups of up to 1.46x on visually-grounded tasks. MASSV provides a scalable, architecture-compatible method for accelerating both current and future VLMs.
@@ 1 Introduction
While SD has been well-studied for text-only models, extending it to vision-language models (VLMs) introduces unique challenges. VLMs process multimodal inputs by mapping image features and text tokens into a joint embedding space, enabling sophisticated visual reasoning capabilities (Radford et al., 2021; Liu et al., 2023). This multimodal conditioning presents two fundamental challenges for SD: (1) architectural incompatibility, as small language models lack the components to process visual inputs, and (2) distribution mismatch, as unimodal draft models cannot effectively capture the visually-grounded nature of the target VLM’s outputs.
Previous approaches have addressed these challenges either by excluding image tokens entirely or by training small multimodal models from scratch (Gagrani et al., 2024). The former approach fails to leverage visual information, while the latter requires substantial computational resources and may still suffer from distribution misalignment. Lee et al. (2024) explored ensemble-based methods that combine multiple drafting strategies through batch inference, achieving robustness across diverse input scenarios without requiring additional model parameters. However, these ensemble approaches do not fundamentally address the distribution mismatch between draft and target models, instead relying on averaging predictions from multiple unaligned drafters. Neither of these approaches fully exploit the potential of existing model families or directly optimize for the distribution alignment needed for effective SD.
@@ 2.1 Speculative Decoding
Formally, let $X_{1:t}=X_{1},X_{2},...,X_{t}$ be the input sequence for the current iteration. $M_{q}$ first autoregressively samples $\gamma$ draft tokens $X_{t+1:t+\gamma}$, where token $X_{t+i}$ is sampled with probability $q(X_{t+i}|X_{1:t+i-1})$. Next, $M_{p}$ computes the probabilities $p(X_{t+i}|X_{1:t+i-1})$ for $i=1,2,...,\gamma+1$ in parallel with one forward call. These probabilities are used to evaluate the draft tokens sequentially, with the probability of accepting token $X_{t+i}$ being $\min\left(1,\frac{p(X_{t+i}|X_{1:t+i-1})}{q(X_{t+i}|X_{1:t+i-1})}\right)$. If the token is accepted, it is added to the generation output and the next token is evaluated. Otherwise, if the token is rejected, a new token is sampled from the residual distribution $\operatorname{norm}(\max(p(\cdot|X_{1:t+i-1})-q(\cdot|X_{1:t+i-1}),0))$ and the iteration ends. Sampling from the residual distribution ensures the output distribution of the speculative decoding algorithm is the same as the target’s output distribution. In the degenerate case where sampling is disabled (temperature = 0), the algorithm simplifies to greedy decoding. The draft model generates tokens by selecting $X_{t+i}=\arg\max_{x}q(x|X_{1:t+i-1})$. During verification, token $X_{t+i}$ is accepted if and only if $X_{t+i}=\arg\max_{x}p(x|X_{1:t+i-1})$. If rejected, the token is set to $\arg\max_{x}p(x|X_{1:t+i-1})$.
@@ 2.2 Vision-Language Models
Vision-language models (VLMs) process multimodal inputs, consisting of visual and text tokens, by mapping the tokens into a joint embedding space. A VLM consists of three components: a vision encoder $\phi_{I}$, multimodal projector $g_{\theta}$, and a language model $M_{p}$. Given an input consisting of tokens $X_{1:t}$ and visual information $I$, a VLM first extracts $m$ features $H_{1:m}=\phi_{I}(I)$ from the image using the vision encoder. These image features are then projected into the joint embeddings space $V_{i}=g_{\theta}(H_{i})$ for $i\in\{1,...,m\}$. Finally, the VLM samples the next token $X_{t+1}$ from $p(\cdot|X_{1:t},V_{1:m})$, where $p(\cdot|\cdot)$ denotes the conditional probability distribution of $M_{p}$. Note that directly using SD to accelerate a VLM on multimodal inputs requires the drafter to also be a VLM. However,  Gagrani et al. (2024) show that a small language model (SLM) can be used as an effective drafter by conditioning it only on the text tokens in the input. Concretely, given an SLM drafter $M_{q}$, the draft token $X_{t+i}$ is sampled from $q(\cdot|X_{1:t+i-1})$ for $i=1,...,\gamma$. We refer to this as text-only drafting and use it as the baseline in our experiments.
@@ 3 Methodology
We introduce a method to adapt an SLM into an effective draft model for LLaVA-style vision–language models, which employ a modular architecture of separate vision encoder and language model components connected via a projection layer that maps image features into the language model’s embedding space. Our approach integrates the target VLM’s frozen vision encoder into the SLM through a randomly initialized MLP-based projector, preserving architectural compatibility while enabling visual processing. We then align the adapted model with the target VLM through a two-phase training protocol: (1) the projector is pretrained on paired image–text data to establish robust visual grounding; and (2) the model undergoes self-distilled visual instruction tuning to optimize token-level distribution alignment.
@@ 3.2 Multimodal Alignment
where $V_{i}=g_{\psi}^{q}(\phi_{I}^{p}(I_{i}))$ are the projected visual features, $y_{i}^{\prime k}$ is the $k$-th token of the target’s response, and $q_{\theta}$ denotes the drafter’s distribution with parameters $\theta=\{\psi,\theta_{q}\}$ (projector and SLM parameters). In contrast to generic visual instruction tuning with fixed dataset labels, our self-distillation strategy trains the drafter on the target’s actual outputs, directly optimizing for the acceptance mechanism in SD. SDViT addresses this through diverse sampling, where the target VLM generates responses across different temperature values with top-p sampling, creating a varied dataset that better represents the full response distribution. Specifically, draft tokens are accepted with probability $\min\left(1,\frac{p(X_{t}|X_{1:t},I)}{q(X_{t}|X_{1:t},I)}\right)$. By training on the target’s outputs rather than generic labels, we maximize the overlap between the drafter’s distribution $q$ and the target’s distribution $p$, leading to higher token acceptance rates during inference. Our experimental results in Section 5.1 show that this alignment translates to improved token acceptance rates during SD.
@@ 4.2 Results
Our results demonstrate MASSV’s significant improvements over the text-only baseline across all evaluated settings (see Table 1). At temperature $T=0$, MASSV achieves a noticeable increase in mean accepted length (MAL), most notably improving by $30.1\%$ (from 2.46 to 3.20) for the Qwen2.5-VL 7B Instruct model. Similarly, at $T=1$, MASSV attains a MAL improvement of $23.3\%$ (from 2.58 to 3.18). These improvements are consistent across different downstream tasks, with the largest relative gains observed in visually intensive tasks such as COCO captioning. For instance, MASSV increases MAL by $47.5\%$ (2.21 to 3.26) on COCO captioning tasks at $T=0$, highlighting the importance of multimodal drafting for visually-grounded generations. Moreover, MASSV consistently outperforms the baseline on the Gemma3 family despite their significant architectural differences (e.g., dynamic visual token count in Qwen2.5-VL versus interleaved sliding window attention in Gemma3). Specifically, MASSV improves MAL by $15.6\%$ (2.76 to 3.19) on Gemma3-12B IT at $T=0$, demonstrating its versatility and effectiveness across diverse VLMs.
@@ 5.1 Effect of Self-Distilled Visual Instruction Tuning
We assess the role of self-distilled distillation in our method by comparing drafters trained with SDViT versus standard fine-tuning on a vanilla dataset. Specifically, we adapt Qwen2.5-1.5B Instruct and Gemma3-1B IT into drafters for Qwen2.5-VL 7B Instruct and Gemma3-12B IT, respectively. Table 2 presents the quantitative evaluation results, clearly demonstrating that self-distilled visual instruction (SDViT) is crucial for our method’s effectiveness. Without SDViT (i.e., MASSVw/o ${}_{\text{w/o SDViT}}$), the Gemma3-1B IT draft model shows a significant performance regression, achieving a mean accepted length of 2.33 compared to the baseline’s 2.74 (i.e., a 13% decrease in mean accepted length). This indicates that naive architectural adaptation without distribution alignment can be notably detrimental to performance. In contrast, when enhanced with SDViT, the model achieves a mean accepted length of 3.14, representing a substantial 14.6% improvement over the baseline and a 1.18× speedup. These results highlight the critical role of distribution alignment in multimodal speculative decoding.
@@
where $P$ and $Q$ are the target and drafter token distributions, respectively, and $\mathcal{X}$ is the vocabulary. TVD is particularly relevant in the context of speculative decoding, as it bounds the expected probability that tokens proposed by the draft model will be rejected by the target model. By minimizing TVD through our SDViT approach, we directly optimize for higher token acceptance rates, which explains the improved mean accepted length observed in our experiments. For discrete distributions like token probabilities, TVD ranges from 0 (identical distributions) to 1 (completely disjoint distributions). Figure 3 shows the resulting distribution.
Figure 3: Histogram of total variation distances (TVD), comparing the Qwen2.5-1.5B drafters trained with (purple) and without (orange) self-distilled visional instruction (SDViT) against the Qwen2.5-VL 7B target model on our multimodal speculative decoding benchmark. MASSV yields a highly skewed distribution concentrated at low TVD values, indicating tighter alignment with the target’s token distribution. In contrast, MASSVw/o ${}_{\text{w/o SDViT}}$ produces a broader, heavier-tailed distribution, reflecting reduced alignment. The left-skewed shape of the MASSV distribution quantitatively suggests that SDViT narrows the distributional gap between draft and target.
@@ 5.2 Text-Only vs Multimodal Drafting
Based on these observations, we focus exclusively on multimodal drafting in our main experiments (Section 4). This choice ensures we capture the full benefits of visual information while maintaining strong distribution alignment through self-distilled visual instruction tuning. As we demonstrate across multiple model families and tasks, this combination of multimodal capability and distribution alignment yields consistent improvements in speculative decoding performance.

### v2

@@ Abstract
Speculative decoding significantly accelerates language model inference by enabling a lightweight draft model to propose multiple tokens that a larger target model verifies simultaneously. However, applying this technique to vision-language models (VLMs) presents two fundamental challenges: small language models that could serve as efficient drafters lack the architectural components to process visual inputs, and their token predictions fail to match those of VLM target models that consider visual context. We introduce Multimodal Adaptation and Self-Data Distillation for Speculative Decoding of Vision-Language Models (MASSV), which transforms existing small language models into effective multimodal drafters through a two-phase approach. MASSV first connects the target VLM’s vision encoder to the draft model via a lightweight trainable projector, then applies self-distilled visual instruction tuning using responses generated by the target VLM to align token predictions. Comprehensive experiments across the Qwen2.5-VL and Gemma3 model families demonstrate that MASSV increases accepted length by up to 30% and delivers end-to-end inference speedups of up to 1.46x compared to conventional text-only drafting baselines on visually-grounded tasks. MASSV provides a scalable, architecture-compatible method for accelerating both current and future VLMs.
@@ 1 Introduction
While SD has been well-studied for text-only models, extending it to vision-language models (VLMs) introduces unique challenges. VLMs process multimodal inputs by mapping image features and text tokens into a joint embedding space, enabling sophisticated visual reasoning capabilities (Radford et al., 2021; Liu et al., 2023). This multimodal conditioning presents two fundamental challenges for SD: (1) architectural incompatibility, as small language models lack the components to process visual inputs, and (2) distribution mismatch, as unimodal draft models cannot effectively capture the visually-grounded nature of the target VLM’s outputs. Previous approaches have addressed these challenges either by excluding image tokens entirely or by training small multimodal models from scratch (Gagrani et al., 2024). The former approach fails to leverage visual information, while the latter requires substantial computational resources and may still suffer from distribution misalignment. Lee et al. (2024) explored ensemble-based methods that combine multiple drafting strategies through batch inference, achieving robustness across diverse input scenarios without requiring additional model parameters. However, these ensemble approaches do not fundamentally address the distribution mismatch between draft and target models, instead relying on averaging predictions from multiple unaligned drafters. Neither of these approaches fully exploit the potential of existing model families or directly optimize for the distribution alignment needed for effective SD.
@@ 2.1 Speculative Decoding
Formally, let $X_{1:t}=X_{1},X_{2},...,X_{t}$ be the input sequence for the current iteration. $M_{q}$ first autoregressively samples $\gamma$ draft tokens $X_{t+1:t+\gamma}$, where token $X_{t+i}$ is sampled with probability $q(X_{t+i}|X_{1:t+i-1})$. Next, $M_{p}$ computes the probabilities $p(X_{t+i}|X_{1:t+i-1})$ for $i=1,2,...,\gamma+1$ in parallel with one forward call. These probabilities are used to evaluate the draft tokens sequentially, with the probability of accepting token $X_{t+i}$ being $\min\left(1,\frac{p(X_{t+i}|X_{1:t+i-1})}{q(X_{t+i}|X_{1:t+i-1})}\right)$. If the token is accepted, it is added to the generation output and the next token is evaluated. Otherwise, if the token is rejected, a new token is sampled from the residual distribution $\operatorname{norm}(\max(p(\cdot|X_{1:t+i-1})-q(\cdot|X_{1:t+i-1}),0))$ and the iteration ends. Sampling from the residual distribution ensures the output distribution of the speculative decoding algorithm is the same as the target’s output distribution.
In the degenerate case where sampling is disabled (temperature = 0), the algorithm simplifies to greedy decoding. The draft model generates tokens by selecting $X_{t+i}=\arg\max_{x}q(x|X_{1:t+i-1})$. During verification, token $X_{t+i}$ is accepted if and only if $X_{t+i}=\arg\max_{x}p(x|X_{1:t+i-1})$. If rejected, the token is set to $\arg\max_{x}p(x|X_{1:t+i-1})$.
@@ 2.2 Vision-Language Models
Vision-language models (VLMs) process multimodal inputs, consisting of visual and text tokens, by mapping the tokens into a joint embedding space. A VLM consists of three components: a vision encoder $\phi_{I}$, multimodal projector $g_{\theta}$, and a language model $M_{p}$. Given an input consisting of tokens $X_{1:t}$ and visual information $I$, a VLM first extracts $m$ features $H_{1:m}=\phi_{I}(I)$ from the image using the vision encoder. These image features are then projected into the joint embeddings space $V_{i}=g_{\theta}(H_{i})$ for $i\in\{1,...,m\}$. Finally, the VLM samples the next token $X_{t+1}$ from $p(\cdot|X_{1:t},V_{1:m})$, where $p(\cdot|\cdot)$ denotes the conditional probability distribution of $M_{p}$.
Note that directly using SD to accelerate a VLM on multimodal inputs requires the drafter to also be a VLM. However,  Gagrani et al. (2024) show that a small language model (SLM) can be used as an effective drafter by conditioning it only on the text tokens in the input. Concretely, given an SLM drafter $M_{q}$, the draft token $X_{t+i}$ is sampled from $q(\cdot|X_{1:t+i-1})$ for $i=1,...,\gamma$. We refer to this as text-only drafting and use it as the baseline in our experiments.
@@ 3 Methodology
We introduce a method to adapt an SLM into an effective draft model for LLaVA-style vision-language models, which employ a modular architecture of separate vision encoder and language model components connected via a projection layer that maps image features into the language model’s embedding space. Our approach integrates the target VLM’s frozen vision encoder into the SLM through a randomly initialized MLP-based projector, preserving architectural compatibility while enabling visual processing. We then align the adapted model with the target VLM through a two-phase training protocol: (1) the projector is pretrained on paired image-text data to establish robust visual grounding; and (2) the model undergoes self-distilled visual instruction tuning to optimize token-level distribution alignment. The overall architecture is illustrated in Figure 2.
@@ 3.2 Multimodal Alignment
where $V_{i}=g_{\psi}^{q}(\phi_{I}^{p}(I_{i}))$ are the projected visual features, $y_{i}^{\prime k}$ is the $k$-th token of the target’s response, and $q_{\theta}$ denotes the drafter’s distribution with parameters $\theta=\{\psi,\theta_{q}\}$ (projector and SLM parameters).
In contrast to generic visual instruction tuning with fixed dataset labels, our self-distillation strategy trains the drafter on the target’s actual outputs, directly optimizing for the acceptance mechanism in SD. SDViT addresses this through diverse sampling, where the target VLM generates responses across different temperature values with top-p sampling, creating a varied dataset that better represents the full response distribution. Specifically, draft tokens are accepted with probability $\min\left(1,\frac{p(X_{t}|X_{1:t},I)}{q(X_{t}|X_{1:t},I)}\right)$. By training on the target’s outputs rather than generic labels, we maximize the overlap between the drafter’s distribution $q$ and the target’s distribution $p$, leading to higher token acceptance rates during inference. Our experimental results in Section 5.1 show that this alignment translates to improved token acceptance rates during SD.
@@ 4.2 Results
Our results demonstrate MASSV’s significant improvements over the text-only baseline across all evaluated settings (see Table 1). At temperature $T=0$, MASSV achieves a noticeable increase in mean accepted length (MAL), most notably improving by $30.1\%$ (from 2.46 to 3.20) for the Qwen2.5-VL 7B Instruct model. Similarly, at $T=1$, MASSV attains a MAL improvement of $23.3\%$ (from 2.58 to 3.18). These improvements are consistent across different downstream tasks, with the largest relative gains observed in visually intensive tasks such as COCO captioning. For instance, MASSV increases MAL by $47.5\%$ (2.21 to 3.26) on COCO captioning tasks at $T=0$, highlighting the importance of multimodal drafting for visually-grounded generations. Moreover, MASSV consistently outperforms the baseline on the Gemma3 family despite their significant architectural differences (e.g., dynamic visual token count in Qwen2.5-VL versus interleaved sliding window attention in Gemma3). Specifically, MASSV improves MAL by $15.6\%$ (2.76 to 3.19) on Gemma3-12B IT at $T=0$, demonstrating its effectiveness across diverse VLMs.
@@ 5.1 Effect of Self-Distilled Visual Instruction Tuning
We assess the role of self-distilled distillation in our method by comparing drafters trained with SDViT versus standard fine-tuning on a vanilla dataset. Specifically, we adapt Qwen2.5-1.5B Instruct and Gemma3-1B IT into drafters for Qwen2.5-VL 7B Instruct and Gemma3-12B IT, respectively. Figure 3 demonstrates the efficacy of MASSV with SDViT (green bar) for Qwen2.5-VL 7B Instruct across diverse multimodal benchmarks. MASSV exhibits substantial performance gains, most prominently in COCO Captioning where the mean accepted length increases from 2.21 to 3.26 tokens (+47.5%). Table 2 summarizes our comprehensive ablation study on SDViT across both target models: Qwen2.5-VL 7B Instruct and Gemma3-12B IT. The quantitative evaluation results clearly demonstrate the critical importance of self-distilled visual instruction tuning for effective multimodal SD. For the Gemma3 architecture, without SDViT (denoted as MASSVw/o ${}_{\text{w/o SDViT}}$), the Gemma3-1B IT draft model exhibits a significant performance regression, with mean accepted length deteriorating to 2.33 compared to the baseline’s 2.74 (a 13% decrease in acceptance rate). This indicates that naive architectural adaptation without distribution alignment can be notably detrimental to performance. In contrast, when enhanced with SDViT, the model achieves a mean accepted length of 3.14, representing a substantial 14.6% improvement over the baseline and a 1.18x speedup. These results highlight the critical role of distribution alignment in multimodal SD.
@@
where $P$ and $Q$ are the target and drafter token distributions, respectively, and $\mathcal{X}$ is the vocabulary. TVD is particularly relevant in the context of SD, as it bounds the expected probability that tokens proposed by the draft model will be rejected by the target model. By minimizing TVD through our SDViT approach, we directly optimize for higher token acceptance rates, which explains the improved mean accepted length observed in our experiments. For discrete distributions like token probabilities, TVD ranges from 0 (identical distributions) to 1 (completely disjoint distributions). Figure 4 shows the resulting distribution.
Figure 4: Histogram of total variation distances (TVD), comparing the Qwen2.5-1.5B drafters trained with (purple) and without (orange) self-distilled visional instruction (SDViT) against the Qwen2.5-VL 7B target model on our multimodal SD benchmark. MASSV yields a highly skewed distribution concentrated at low TVD values, indicating tighter alignment with the target’s token distribution. In contrast, MASSVw/o ${}_{\text{w/o SDViT}}$ produces a broader, heavier-tailed distribution, reflecting reduced alignment. The left-skewed shape of the MASSV distribution quantitatively suggests that SDViT narrows the distributional gap between draft and target.
@@ 5.2 Text-Only vs Multimodal Drafting
Based on these observations, we focus exclusively on multimodal drafting in our main experiments (Section 4). This choice ensures we capture the full benefits of visual information while maintaining strong distribution alignment through SDViT. As we demonstrate across multiple model families and tasks, this combination of multimodal capability and distribution alignment yields consistent improvements in SD performance.
@@ 7 Conclusion
8 Acknowledgements
We would like to express our sincere gratitude to Eric Sather and Mike Lasby for their thoughtful feedback and fruitful discussions that helped improve the final manuscript. We also acknowledge Rohan Gupta and Eugene Osovetsky for their valuable insights which emphasized the critical importance of ML software/hardware co-design in addressing these complex challenges.